A 

66−L 

sample 

of 

dry 

air 

cools 

from 

101 

°C 

to 

−25 

°C 

while 

the 

pressure 

...

A 

66−L 

sample 

of 

dry 

air 

cools 

from 

101 

°C 

to 

−25 

°C 

while 

the 

pressure 

is 

maintained 

at 

2.71 atm. What is the final volume?
Answer #1

At constant pressure an constant number of moles, V is directly proportional to Temperature

V1/T1 = V2/T2

66/374 = V2/248

final volume V2 = 43.7647L